OpenStudy (anonymous):

a line having a slope of 3/4 passes through the point (-8,4).write the equation of this line in slope intercept form

OpenStudy (anonymous):

Use the fomula y-y1=m(x-x1)

OpenStudy (anonymous):

it's called the point slope fomula if you want to look it up

OpenStudy (anonymous):

equation of a line in point-slope form is\[y-y _{1}=m(x-x _{1})\] where m is the slope of the line. (x1,y1) is the point which its passing through.. now, plugging in the values you get, \[y-4=(3/4)(x+8)\]\[=> 4y-16=3x+21 => 4y-3x=21+16 => 4y-3x=37\]

OpenStudy (amistre64):

y = (3/4)x + B ; when (x=-8, y=4) 4 = 3(-8)/4 +B 4 = 3(-2) + B 10 = B y = (3/4)x + 10